# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v2

@@ Abstract
Because it is difficult to precisely specify complex objectives, reinforcement learning policies are often optimized using flawed proxy rewards that seem to capture the true objective. However, optimizing proxy rewards frequently leads to reward hacking: the optimized reward function ceases to be a good proxy, and the resulting policy performs poorly with respect to the unspecified true reward. Principled solutions to reward hacking have been impeded by the lack of a good definition for the problem. To address this, we introduce a definition of reward hacking based on the correlation between proxy and true rewards for states and actions seen by a “base policy” that breaks down under optimization. We show that this definition captures reward hacking behavior across several realistic settings, including in reinforcement learning from human feedback (RLHF). We then show theoretically that regularization to the base policy can effectively prevent reward hacking. While current RLHF approaches apply a KL penalty between the action distributions of policies, our theory suggests that it is more effective to regularize using the $\chi^{2}$ divergence between the policies’ occupancy measures. We intuitively show why this type of regularization is superior and demonstrate that it better mitigates reward hacking in practice across four realistic domains, including RLHF for LLMs. Our code is available at https://github.com/cassidylaidlaw/orpo.
@@ 1 Introduction
A major challenge for the designers of goal-oriented AI systems is specifying a reward function that robustly captures their goals and values. Manually designing reward functions is difficult due to the ambiguities and complexity underlying real-world scenarios (Ibarz et al., 2018). An alternative is to learn reward functions from human data (Sadigh et al., 2017; Jeon et al., 2020), but these often fail to generalize outside the distribution of behavior seen during training (McKinney et al., 2023; Tien et al., 2023). Thus, a learned or hand-specified reward function is often just a proxy for the true reward underlying the system designer’s intent. Misalignment between the two objectives can lead to reward hacking: a learned policy performs well according to the proxy reward function, but not according to the true reward function (Russell et al., 2010; Amodei et al., 2016; Pan et al., 2022; Skalse et al., 2022). A reward hacking policy’s behavior is often undesirable and can be especially catastrophic when deployed in safety-critical scenarios, such as autonomous driving (Krakovna et al., 2019; Turner et al., 2019; Knox et al., 2022). Unfortunately, reward hacking is a common phenomenon (Krakovna, 2018) and has harmful effects in the real world (Lum & Isaac, 2016; Corbett-Davies et al., 2017; Obermeyer et al., 2019; Milli et al., 2021; Franchi et al., 2023; Kleinberg et al., 2023).
The ideal solution to prevent reward hacking would be to perfectly align the specified proxy and unknown true reward; however, in many domains, this is impossible to achieve. For example, imagine trying to design a reward function for a self-driving car. It would have to capture the speed at which the car reached the destination, comfort of the passenger, all applicable laws, and other factors, many of which are difficult to robustly measure; furthermore, these factors would have to be carefully weighted against each other (Knox et al., 2022). In practice, reward hacking can occur even with significant reward engineering efforts. For example, hand-designed reward functions for recommender systems have led to adverse outcomes in terms of user experience (Stray et al., 2022). Reinforcement learning from human feedback (RLHF), where a learned reward function is used to align a large-language model (LLM) based assistant to human preferences, provides another example. Without regularization, RLHF leads to the overoptimization of the learned reward function, a case of reward hacking (Gao et al., 2022).
Since proxy reward functions for complex tasks are nearly always misspecified in practice, what can be done to avoid reward hacking? There is a lack of principled solutions for preventing reward hacking, stemming from a more fundamental problem: defining reward hacking in a formal sense that captures realistic cases. In particular, it is difficult to characterize what makes a proxy reward function “good”. For example, Skalse et al. (2022) define a notion of “unhackability” for proxy reward functions but find that it can only hold if the proxy is a multiple of the true reward, which is obviously unrealistic.
@@
We argue that proxies are chosen because they seem to capture the true objective under some base distribution of behavior. To formalize this, we define a proxy reward as one that correlates with the true reward function under the distribution of states and actions visited by a base policy. For example, if we use a human-like policy as the base policy for self-driving cars, then a proxy reward would have to correlate with the true reward for states and actions seen during typical human driving. In the case of RLHF, where the reward function is learned from human preference feedback over different model responses, a natural base policy is the supervised fine-tuning (SFT) model used to generate those responses. If the reward learning process is effective, then the learned reward should correlate well with the true human-judged reward on the distribution of responses it was trained on, making it a good proxy.
We thus formally define reward hacking as occurring when optimizing a proxy reward induces a decrease in true reward compared to the base policy (Figure 1). We show that this definition and our definition of a proxy reward function capture a number of intuitive cases of reward hacking in realistic environments, including traffic control, blood glucose regulation, and RLHF (Figure 2).
Furthermore, we show that our definition allows us to derive a method for avoiding reward hacking during policy optimization by regularizing the optimized policy to be similar to the base policy. Specifically, we find that optimizing the proxy reward minus a regularization term provides a provable lower bound on improvement in the true reward. The amount of regularization needed depends on how strong the correlation is between the proxy and true rewards; as the correlation becomes stronger, it is possible to use less regularization.
@@
Regularizing policy optimization to a base policy is already used in practice during RLHF via a KL divergence penalty (Stiennon et al., 2020; Bai et al., 2022), and Proposition 5.1 provides some theoretical justification for why this works. However, while the regularization in RLHF penalizes the KL divergence between the optimized and base policies’ action distributions, our result suggests that it is better to penalize the chi-squared ($\chi^{2}$) divergence between the policies’ occupancy measures. A policy’s occupancy measure (OM) is the distribution of states or state-action pairs seen by the policy when it interacts with its environment. Unlike action distribution (AD)-based metrics, occupancy measures take into account the states that the agent reaches, not just the actions it takes. While algorithms based on occupancy measures have been widely used for imitation learning (Ho & Ermon, 2016), offline RL (Lee et al., 2022), and efficient exploration (Hazan et al., 2019), using OM divergences to prevent reward hacking remains unexplored. We compare OM vs. AD regularization and $\chi^{2}$ vs. KL divergence, presenting intuitive reasons why $\chi^{2}$ OM divergence may be a better regularization target in Figures 3 and 4.
@@
Based on our theoretical results, we empirically investigate using the benefits of using occupancy measure regularization to prevent reward hacking. We implement OM-based regularization in practice using a discriminator network that approximates the OM divergence between policies. We then optimize policies with misaligned proxy reward functions in multiple reward hacking benchmark environments (Pan et al., 2022) using OM and AD regularization. The results of our experiments demonstrate that training with occupancy measure regularization leads to better performance under the unseen true reward function in all of the environments, validating our theoretical results. In contrast, we find that it is difficult to tune AD regularization in some environments to both prevent reward hacking and allow meaningful improvement over the base policy. Furthermore, we compare regularizing with $\chi^{2}$ divergence, which our theoretical results suggest using, to KL divergence, the prior standard in settings like RLHF. We find that regularization with chi-squared divergence leads to more stable results across regularization coefficients and often achieves higher true reward.
@@
    Using our definition, we establish that optimizing a proxy reward with $\chi^{2}$ occupancy measure regularization leads to a provable improvement in the unknown true reward function.
@@
    We implement $\chi^{2}$ OM regularization in practice and demonstrate that it is superior to the current standard for preventing reward hacking via regularization.
@@ 2 Related work
While there have been separate lines of work investigating reward hacking and exploring the use of occupancy measure divergences for other applications, to the best of our knowledge, we are the first to specifically study applying occupancy measure regularization to the problem of reward hacking.
Reward hacking.  Some prior works establish theoretical models of reward hacking as a special case of Goodhart’s Law (Goodhart, 1984; Leike et al., 2018; Manheim & Garrabrant, 2019; Krakovna, 2019; Skalse et al., 2022; Ngo et al., 2023). Krakovna (2018) provide a list of many examples of reward hacking. Pan et al. (2022) categorize types of reward misspecification and relate optimization power to reward hacking.
Safe reinforcement learning.  Regularizing policies to be similar to an offline policy based on their action distribution KL divergence was proposed by Stiennon et al. (2020) and has since been widely employed in the context of optimizing LLMs using RLHF (Ouyang et al., 2022; Bai et al., 2022; Glaese et al., 2022). KL regularization for RLHF has been further studied by Vieillard et al. (2021), Gao et al. (2022), and Korbak et al. (2022). Some alternative approaches to avoid reward hacking include quantilizers (Taylor, 2016), “mild” optimization (Taylor et al., 2020), and impact regularization (Turner et al., 2020). While constrained RL can prevent the misbehavior of agents that optimize flawed reward functions (Dalal et al., 2018; Chow et al., 2019; Zhang et al., 2020; Roy et al., 2022), it simply shifts the difficulty of designing a reward function to specifying a set of constraints and weights. Robust RL usually considers a misspecified transition model, but some work has explored misspecified reward functions (Derman et al., 2021; Gadot et al., 2024). Other proposals to address the reward specification problem attempt to infer the true reward function based on the given proxy reward function, environment context, and/or feedback from humans (Hadfield-Menell et al., 2017; Reddy et al., 2020; Lee et al., 2021).
@@
Offline reinforcement learning.  Prior work in offline RL may appear to be particularly related to our work. Many offline RL algorithms use occupancy measure or action distribution-based regularization to ensure that the learned policy remains within the training data distribution (Lee et al., 2022; Mandal et al., 2023; He, 2023; Cheng et al., 2022; Rashidinejad et al., 2023; Xie et al., 2023). However, the settings are fundamentally different: while offline RL is limited by a lack of coverage in the training data, the difficulty in our setting is that the reward function is misspecified. While it might be possible to avoid reward hacking by using offline RL algorithms, we leave this to future work and focus on regularization-based approaches in the online RL setting.
@@
Figure 2: We show that reward hacking in four realistic environments and an illustrative gridworld is captured by our definition. The top row shows the distribution of proxy and true reward values for state-action pairs sampled from a natural base policy for each environment; in all environments, the proxy and true rewards are correlated. However, if the proxy is optimized via reinforcement learning, the correlation breaks down and the true reward is lower than the base policy, which we define as reward hacking (middle row). Our theoretical results show that RL with occupancy measure regularization to the base policy can prevent reward hacking and enable an increase in true reward (bottom row).
@@ 3 Preliminaries
To study reward hacking, we consider the setting of an infinite-horizon Markov decision process (MDP). An agent takes actions $a\in\mathcal{A}$ to transition between states $s\in\mathcal{S}$ over a series of timesteps $t=0,1,2,\ldots$. We assume that $\mathcal{S}$ and $\mathcal{A}$ are finite for simplicity but our results can easily generalize to infinite state or action spaces. The first state $s_{0}$ is sampled from an initial distribution $\mu_{0}(s)$, and when an agent takes action $a_{t}$} in $s_{t}$} at time t{t}, the next state $s_{t+1}$ is reached at timestep $t+1$ with transition probability $p(s_{t+1}\mid s_{t},a_{t})$. The agent aims to optimize a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and rewards are accumulated over time with discount factor $\gamma\in[0,1)$. A policy $\pi$ maps each state $s$ to a distribution over actions to take at that state $\pi(a\mid s)$. We define the (normalized) return of a policy $\pi$ under a reward function $R$ as
@@
Intuitively, the occupancy measure represents the distribution of states and actions visited by the policy over time. Furthermore, combining (1) and (2), it is easy to show that the return of a policy can be re-written as the expected reward for states and actions sampled from the occupancy measure:
@@
Desiderata for a definition of reward hacking.  To understand why defining reward hacking is difficult, consider the case study of RLHF. In this case, optimizing a learned reward function over LLM outputs eventually leads to the LLM producing nonsensical responses. This clearly satisfies our informal definition of “optimizing the proxy makes the true reward go down.” However, what if we were to optimize $\tilde{R}(s,a)=-R(s,a)$? In this case, optimizing $\tilde{R}$ also makes the true reward go down, but arguably this is not reward hacking because $\tilde{R}$ was not a good “proxy” in the first place. Thus, a good definition of reward hacking should distinguish between “reasonable” proxies and reward functions that are clearly unrelated (or opposite) to the true reward function.
In the RLHF example, optimizing until the LLM produces nonsensical outputs seems like bad enough behavior that we would say the proxy has been “hacked.” However, what if optimizing the proxy produced mediocre but not obviously bad outputs? Even if optimizing the proxy does not lead to near-optimal true reward, we would not always say that reward hacking has occurred. Thus, a good definition of reward hacking should also choose a threshold for true reward: below the threshold, reward hacking is occurring, and above the threshold it is not.
Prior definitions of reward hacking.  Because of these difficulties, prior work has struggled to define reward hacking. With regard to our first desideratum, defining reasonable proxies, Skalse et al. (2022) introduce a notion of an “unhackable” proxy reward but find that a proxy is only unhackable if it is equivalent to the true reward up to scaling; this is clearly too restrictive of a definition. The robust MDP literature has considered proxy rewards that differ from the true reward by at most some constant, e.g., $|R(s,a)-\tilde{R}(s,a)|\leq c$ for all $s,a$ (Derman et al., 2021). However, we find that in many realistic cases of reward hacking, there may be some states where a proxy differs from the true reward by an arbitrarily large amount, and so these definitions are generally not applicable. With regard to the second desideratum, defining a threshold of true reward, there is little existing literature.

### v3

@@ Abstract
Because it is difficult to precisely specify complex objectives, reinforcement learning policies are often optimized using proxy reward functions that only approximate the true goal. However, optimizing proxy rewards frequently leads to reward hacking: the optimized reward function ceases to be a good proxy and the resulting policy performs poorly with respect to the unspecified true reward. Principled solutions to reward hacking have been impeded by the lack of a good definition for the problem. To address this gap, we introduce a definition of reward hacking based on the correlation between proxy and true rewards for states and actions seen by a “base policy” that breaks down under optimization. We show that this definition captures reward hacking behavior across several realistic settings, including in reinforcement learning from human feedback (RLHF). Using our formulation, we show theoretically that regularization to the base policy can effectively prevent reward hacking. While the current practice in RLHF applies a KL penalty between action distributions for this purpose, our theory suggests regularizing the $\chi^{2}$ divergence between the policies’ occupancy measures can be more effective. We intuitively show the benefits of this type of regularization and demonstrate that it better mitigates reward hacking in practice across four realistic settings, including RLHF. Our code is available at https://github.com/cassidylaidlaw/orpo.
@@ 1 Introduction
A major challenge for the designers of goal-oriented AI systems is specifying a reward function that robustly captures their goals and values. Manually designing reward functions is difficult due to the ambiguities and complexity underlying real-world scenarios (Ibarz et al., 2018). An alternative is to learn reward functions from human data (Sadigh et al., 2017; Jeon et al., 2020), but these often fail to generalize outside the distribution of behavior seen during training (McKinney et al., 2023; Tien et al., 2023). Thus, a learned or hand-specified reward function is often just a proxy for the true reward underlying the system designer’s intent. Misalignment between the two objectives can lead to reward hacking: a learned policy performs well according to the proxy reward function but not according to the true reward function (Russell et al., 2010; Amodei et al., 2016; Pan et al., 2022; Skalse et al., 2022). A reward hacking policy’s behavior is often undesirable and can be especially catastrophic when deployed in safety-critical scenarios, such as autonomous driving (Krakovna et al., 2019; Turner et al., 2019; Knox et al., 2022). Unfortunately, reward hacking is a common phenomenon (Krakovna, 2018) and has harmful effects in the real world (Lum & Isaac, 2016; Corbett-Davies et al., 2017; Obermeyer et al., 2019; Milli et al., 2021; Franchi et al., 2023; Kleinberg et al., 2023).
The ideal solution to prevent reward hacking would be to perfectly align the specified proxy and unknown true reward; however, in many domains, this is impossible to achieve. For example, imagine trying to design a reward function for a self-driving car. It would need to capture multiple factors: the arrival time, passenger comfort, compliance with traffic laws, and various other considerations—many of which are difficult to robustly measure and would need to be carefully weighted against each other (Knox et al., 2022). In practice, reward hacking can occur even with significant reward engineering efforts. This is evident in the case of recommender systems, where hand-designed reward functions have led to adverse outcomes in terms of user experience (Stray et al., 2022).
Since proxy reward functions for complex tasks are nearly always misspecified in practice, what can be done to avoid reward hacking? There is a lack of principled solutions for preventing reward hacking, stemming from a more fundamental problem: defining reward hacking in a formal sense that captures realistic cases. In particular, it is difficult to characterize what makes a proxy reward function “good”. For example, Skalse et al. (2022) define a notion of “unhackability” but find that it can only hold if such an objective is a multiple of the true reward—a requirement that is unlikely in practice.
@@
We argue that proxies are chosen because they seem to capture the true objective under some base distribution of behavior. To formalize this, we define a proxy reward as one that correlates with the true reward function under the distribution of states and actions visited by a base policy. Then, we define a hackable proxy as one that induces a decrease in true reward compared to the base policy when optimized (Figure 1). This framework captures a wide range of intuitive reward hacking cases across realistic environments: traffic control, pandemic mitigation, blood glucose regulation, and reinforcement learning from human feedback (RLHF) (Figure 2).
@@
Furthermore, our definition motivates a method for avoiding reward hacking during policy optimization by regularizing the optimized policy to be similar to the base policy. Specifically, we find that optimizing the proxy reward minus a regularization term provides a provable lower bound on improvement in true reward. The amount of regularization needed depends on the correlation between the proxy and true rewards; as the correlation becomes stronger, it is possible to use less regularization. In practice, RLHF already uses this form of regularization through a KL divergence penalty to the base policy (Stiennon et al., 2020; Bai et al., 2022), and Theorem 5.1 provides theoretical justification for why this works. However, our results suggest an important modification: rather than penalizing the KL divergence between the policies’ action distributions as currently done in RLHF, we should instead penalize the chi-squared ($\chi^{2}$) divergence between the policies’ occupancy measures. A policy’s occupancy measure (OM) is the distribution of states or state-action pairs seen by the policy when it interacts with its environment. Unlike action distribution (AD)-based metrics, occupancy measures capture not just the actions an agent takes, but also the states it reaches. We compare both OM to AD regularization and $\chi^{2}$ to KL divergence, presenting intuitive reasons why $\chi^{2}$ OM divergence may be a better regularization target in Figures 3 and 4.
@@
Based on our theoretical results, we empirically investigate the benefits of using $\chi^{2}$ occupancy measure regularization to prevent reward hacking. Following previous work (Ho & Ermon, 2016), we implement OM regularization using a discriminator network that approximates the OM divergence between policies. We evaluate our approach in multiple reward hacking benchmark environments (Pan et al., 2022), comparing the effects of OM and AD regularization with both $\chi^{2}$ and KL divergence on policies optimized with hackable proxy reward functions. The results of our experiments demonstrate that training with occupancy measure regularization leads to better performance under the unseen true reward function in all of the environments, validating our theoretical results. In contrast, we find that it is difficult to tune AD regularization in some environments to both prevent reward hacking and allow meaningful improvement over the base policy. Furthermore, regularization with $\chi^{2}$ divergence tends to be more robust to the choice of regularization coefficient and often achieves higher true reward than regularizing with KL divergence.
@@
    Using our definition, we establish that optimizing a proxy reward with $\chi^{2}$ occupancy measure (OM) regularization leads to a provable improvement in the unknown true reward function.
@@
    We show how to practically implement $\chi^{2}$ OM regularization and demonstrate that it outperforms the current standard for preventing reward hacking via action distribution (AD) regularization.
@@ 2 Related Work
While there have been separate lines of work investigating reward hacking and exploring the use of occupancy measure divergences for other applications, to the best of our knowledge, we are the first to specifically study the application of occupancy measure regularization to the problem of reward hacking.
Reward hacking.  Some prior works establish theoretical models of reward hacking as a special case of Goodhart’s Law (Goodhart, 1984; Leike et al., 2018; Manheim & Garrabrant, 2019; Krakovna, 2019; Skalse et al., 2022; Ngo et al., 2023). Krakovna (2018) catalog many examples of reward hacking. Pan et al. (2022) categorize different types of reward misspecification and relate optimization power to reward hacking.
Safe reinforcement learning.  In the context of reward hacking, regularizing policies to be similar to an offline policy based on their action distribution KL divergence was proposed by Stiennon et al. (2020) and has since been widely employed in the context of optimizing LLMs using RLHF (Ouyang et al., 2022; Bai et al., 2022; Glaese et al., 2022). KL regularization for RLHF has been further studied by Vieillard et al. (2021), Gao et al. (2022), and Korbak et al. (2022). Nika et al. (2024) propose a type of occupancy measure regularization in RLHF but it is limited to deterministic MDPs, while we study general stochastic MDPs. Some alternative approaches to avoid reward hacking include quantilizers (Taylor, 2016), “mild” optimization (Taylor et al., 2020), and impact regularization (Turner et al., 2020). While constrained RL can prevent the misbehavior of agents that optimize flawed reward functions (Dalal et al., 2018; Chow et al., 2019; Zhang et al., 2020; Roy et al., 2022), it simply shifts the difficulty of designing a reward function to specifying a set of constraints and weights. Robust RL usually considers a misspecified transition model, but some work has explored misspecified reward functions (Derman et al., 2021; Gadot et al., 2024). Other proposals to address the reward specification problem attempt to infer the true reward function based on the given proxy reward function, environment context, and/or feedback from humans (Hadfield-Menell et al., 2017; Reddy et al., 2020; Lee et al., 2021). Gleave et al. (2021) have previously studied quantifying the similarity of reward functions.
Other divergences for regularization in RLHF.  Some work has explored using divergences other than KL divergence for regularization in RLHF. Go et al. (2023) investigate using other $f$-divergences for aligning language models. Wang et al. (2023a) explore generalizing direct preference optimization (DPO) (Rafailov et al., 2024), a type of RLHF that leverages offline RL, to other $f$-divergences. The most relevant work to ours in this area is Huang et al. (2024), which also proposes using $\chi^{2}$ divergence as a more theoretically grounded alternative to KL divergence for regularization in RLHF. Our work differs from theirs in several key aspects: we address reward hacking in general rather than focussing specifically on learning from preference data; we use online RL instead of their DPO-like algorithm; and, we validate our approach through experiments in realistic environments while their work is primarily theoretical.
@@
Offline reinforcement learning.  Our work bears a superficial resemblance to prior work in offline RL, where algorithms commonly use occupancy measure or action distribution-based regularization to constrain the learned policy within the training data distribution (Fujimoto et al., 2019; Lee et al., 2022; Mandal et al., 2023; He, 2023; Cheng et al., 2022; Rashidinejad et al., 2023; Xie et al., 2023). However, the settings are fundamentally different: while offline RL is limited by a lack of coverage in the training data, the difficulty in our setting is that the reward function is misspecified. While it might be possible to avoid reward hacking by using offline RL algorithms, we leave this to future work and focus on regularization-based approaches in the online RL setting.
Refer to captionSusceptibleExposedInfectiousRecovered
@@
Figure 2: Our definition of reward hacking successfully describes reward hacking behavior in four realistic environments and an illustrative gridworld. The top row shows the distribution of proxy and true reward values for state-action pairs sampled from a domain-appropriate base policy for each environment; in all environments, the proxy and true rewards are correlated. However, as shown in the middle row, this correlation breaks down if the proxy is optimized via RL, and the true reward achieved is lower than that of the base policy, which we characterize as reward hacking. In line with our theoretical results, the bottom row shows that RL with occupancy measure regularization to the base policy can prevent reward hacking while enabling an increase in true reward.
@@ 3 Preliminaries
To study reward hacking, we consider the setting of an infinite-horizon Markov decision process (MDP). An agent takes actions $a\in\mathcal{A}$ to transition between states $s\in\mathcal{S}$ over a series of timesteps $t=0,1,2,\ldots$. We assume that $\mathcal{S}$ and $\mathcal{A}$ are finite for simplicity, but our results can easily generalize to infinite state or action spaces. The first state $s_{0}$ is sampled from an initial distribution $\mu_{0}(s)$, and when an agent takes action $a_{t}$} in $s_{t}$} at time t{t}, the next state $s_{t+1}$ is reached at timestep $t+1$ with transition probability $p(s_{t+1}\mid s_{t},a_{t})$. The agent aims to optimize a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and rewards are accumulated over time with discount factor $\gamma\in[0,1)$. A policy $\pi$ maps each state $s$ to a distribution over actions to take at that state $\pi(a\mid s)$. We define the (normalized) return of a policy $\pi$ under a reward function $R$ as
@@
Intuitively, the occupancy measure represents the distribution of states and actions visited by the policy over time. Furthermore, combining (1) and (2), we can see that the return of a policy can be re-written as the expected reward for states and actions sampled from the occupancy measure:
@@
Desiderata for a definition of reward hacking.  To understand why defining reward hacking is difficult, consider the case study of RLHF. In RLHF, optimizing a learned reward function over LLM outputs eventually leads to the LLM producing nonsensical responses Stiennon et al. (2020). This clearly satisfies our informal definition of “optimizing the proxy makes the true reward decrease.” However, what if we were to optimize $\tilde{R}(s,a)=-R(s,a)$? In this case, optimizing $\tilde{R}$ also makes the true reward go down, but arguably this is not reward hacking because $\tilde{R}$ was not a good “proxy” in the first place. Thus, a good definition of reward hacking should distinguish between “reasonable” proxies and reward functions that are clearly unrelated to the true reward function.
In the example of RLHF, when optimization leads to nonsensical outputs, we can clearly identify that the proxy has been “hacked.” However, what if optimizing the proxy produced mediocre but not obviously bad outputs? Even if optimizing the proxy does not lead to near-optimal true reward, we wouldn’t necessarily classify this behavior as reward hacking. Thus, a good definition of reward hacking should specify a threshold for the true reward: if the true reward falls below the threshold, reward hacking has occurred, and above it, it has not.
Prior definitions of reward hacking.  Due to these challenges, prior work has struggled to define reward hacking. Skalse et al. (2022) attempted to address our first desideratum of defining reasonable proxies by introducing a notion of an “unhackable” proxy reward. However, they found that a proxy is only unhackable if it is equivalent to the true reward up to scaling—a definition that is clearly too restrictive. The robust MDP literature has considered proxy rewards that differ from the true reward by at most some constant, e.g., $|R(s,a)-\tilde{R}(s,a)|\leq c$ for all $s,a$ (Derman et al., 2021). However, we find that in many realistic cases of reward hacking, proxies can differ from the true reward by arbitrarily large amounts in certain states, making these definitions inapplicable. As for our second desideratum of defining a true reward threshold, there is little existing literature.